In a​ year, Marco buys 2 skip trips at ​$ 1500 each and pays ​$ 40 for​ 10,000 gallons of tap water.

Marco is maximizing total utility. If Marco's marginal utility from water is 0.4 unit per​ gallon, his marginal utility from a ski trip is () units.
Answer #1

When utility is maximized, the marginal utility per dollar spent on both the goods equal.

Here, the price paid for a gallon of water = $40/10,000 = $0.004.

Therefore, the marginal utility per dollar spent on water = 0.4/0.004 = 10

So, the marginal utility per dollar spent on skip trips would also be 10.

So, Marginal utility of skip trip/1,500 = 10

So, marginal utility of skip trip = 1,500 * 10 = 15,000.